OpenStudy (anonymous):

Insert grouping symbols to make the following statement true: -8x3-4^2÷2+9x3-2=49

OpenStudy (anonymous):

Oh this looks fun.

OpenStudy (anonymous):

Can the grouping symbols be anywhere

OpenStudy (anonymous):

(-8 x (3 - (4^2/2))) + 9 * (3-2) = 49 I don't know how i did that. but magic. =D

OpenStudy (anonymous):

So it simplifies to (-8 x (3 - (16/2))) + 9 * (3-2) = 49 which is ( -8 x (3 - 8)) + 9 * (1) = 49 which is (-8 x (-5)) + 9 = 49 and finally gets you 40 + 9 = 49!!! Took me a while but that's how you group it

OpenStudy (anonymous):

Thank you !!!! There were several of us trying to get this answer!!! There's another answer someone responded to the same question, but the brackets didn't seem right. THESE DO

OpenStudy (anonymous):

Yeah, this question was interesting... and challenging. Took some brainpower right there.

OpenStudy (anonymous):

That's exactly what intrigued me about it too. Thanks and i'm sure this will be looked up for future reference as i saw it brought up online in 2009!! :) thanks again